# Yi-Fang Wang, Chia-Ming Lee, Yi-Shiuan Chou

### v2

@@ 3.2 Quantitative Results
TABLE II: Quantitative comparison of the noisy face sequences under different masking rations $m_{r}$ between the proposed ML-SELF and other state-of-the-art methods.Highlighting the best performance in red and the second-best performance in blue, considering the utilization of the FF++ Celeb-DF, and DFDC datasets alongside the $m_{r}$ variable.
$m_{r}$ FF++ [1] Celeb-DF [3] DFDC [2]
ACC F1 AUC ACC F1 AUC ACC F1 AUC
Xception [6] 0.0 0.925 0.894 0.972 0.861 0.806 0.910 0.953 0.910 0.981
@@
UCF [30] 0.0 0.937 0.911 0.982 0.856 0.792 0.891 0.890 0.815 0.939
@@
CORE [26] 0.0 0.948 0.925 0.984 0.953 0.940 0.989 0.950 0.903 0.977
0.4 0.883 0.799 0.888 0.858 0.790 0.890 0.907 0.781 0.870
0.8 0.818 0.601 0.663 0.764 0.572 0.661 0.863 0.595 0.651
TALL-Swin [53] 0.0 0.913 0.868 0.881 0.913 0.933 0.924 0.911 0.812 0.984
@@
DFIL [28] 0.0 0.954 0.939 0.987 0.957 0.954 0.964 0.940 0.881 0.955
0.4 0.876 0.808 0.893 0.695 0.684 0.825 0.886 0.720 0.813
0.8 0.759 0.603 0.665 0.518 0.350 0.644 0.855 0.565 0.621
GRACE [Our] 0.0 0.962 0.942 0.989 0.989 0.968 0.998 0.969 0.942 0.988
@@
TABLE III: Comparison of different methods in terms of FLOPs (Floating-point Operations), MACs (Multiply-Accumulate Operations), and number of parameters (##\##Params.).
Method FLOP (T) MACs (T) ##\##Params (M)
@@
DFIL [28] 60.976 30.356 20.811
GRACE (Ours) 70.751 35.246 29.661
@@
The primary performance assessment comparing the handling of invalid facial images between our proposed model, GRACE, and various state-of-the-art schemes is provided in Table II. Under optimal conditions, where most facial images are valid, GRACE exhibits competitive results, holding its own against other cutting-edge DeepFake video detection methods such as Xception [6], MesoNet [5], $F^{3}$-Net [32], RECCE [27], CORE [26], TALL-Swin [53], and DFIL [28]. It is worth noting that TALL-Swin [53] is a video-based approach.
@@
Specifically, the F1-Score of GRACE for DeepFake video detection, when evaluated on FF++ [1], Celeb-DF [3], and DFDC [2], slightly surpasses those of its contemporaries under clean cases (i.e., $m_{r}=0$). This outcome implies that the proposed GRACE with Graph Laplacian is effective and reliable for DeepFake video detection. However, in scenarios where partial face images are invalid due to purposeful attacks on face detectors, the performance of traditional frame-level methods, including Xception [6], MesoNet [5], $F^{3}$-Net [32], RECCE [27], CORE [26], UCF [30], and DFIL [28], may substantially deteriorate since they fail to consider noisy face sequences in real-world scenarios.
@@
Similarly, the video-level DeepFake detection methods, TALL-Swin [53], which heavily rely on temporal cues, may suffer further performance degradation when the masking ratio increases. Invalid faces can cause landmark detection failures and incorrect temporal trajectories. Consequently, the F1-Score of TALL-Swin under a masking ratio of 0.8 in the testing phase is lower than 0.7, implying that all predictions would be categorized as either entirely fake or real. Likewise, the performance of another state-of-the-art video-based DeepFake detection method, TALL-Swin [53], is poor when $m_{r}$ is increased. In stark contrast, all quality indices of our proposed GRACE, evaluated on different datasets, display promising results, suggesting that GRACE is robust and reliable even under highly noisy face sequences (e.g., when $m_{r}=0.8$). Remarkably, since most DeepFake detection methods fail to discuss the impact of unreliable face sequences, the degraded performance is most likely predictable.
@@
To further demonstrate the efficiency and practicality of the proposed GRACE method, we conduct a comprehensive complexity analysis and compare it with other state-of-the-art DeepFake detection methods. Table III presents the comparison results in terms of floating-point operations (FLOPs), multiply-accumulate operations (MACs), and the number of parameters for each method with $16\times 3\times 144\times 144$ tensor for the fair comparison. It is evident that GRACE achieves a remarkable balance between computational complexity and performance. With 70.751 trillion FLOPs, 35.246 trillion MACs, and 29.661 million parameters, GRACE exhibits a moderate computational overhead compared to other methods, such as TALL-Swin [29], UCF [30], and RECCE [27]. Notably, GRACE outperforms these methods in terms of FLOPs and MACs while maintaining a comparable number of parameters. Moreover, GRACE demonstrates superior performance in handling noisy face sequences, as shown in the experimental results, despite having a similar complexity to methods like CORE [26], Xception [6], and DFIL [28]. This highlights the effectiveness of the proposed feature entanglement, graph convolutional network, and graph Laplacian regularization techniques in learning discriminative and robust representations for DeepFake detection. The complexity analysis further substantiates GRACE as a practical and efficient solution for real-world DeepFake detection challenges, offering a compelling trade-off between computational resources and detection accuracy.
@@
Figure 3: The performance comparison of the proposed GRACE and other state-of-the-art methods in terms of AUC under different masking ratios for (a) FF++[1], (b) DFDC [2], and (c) Celeb-DF [3].
The detailed quantitative results, evaluated on the FF++ [1], Celeb-DF [3], and DFDC [2] datasets, are illustrated in Figures 3(a) and 3(c), respectively. In the clean case, i.e., when $m_{r}=0$, the performance of the proposed method is comparable to other state-of-the-art methods. It is observed that performance degradation becomes increasingly pronounced with a rise in the masking ratio during the testing phase, particularly when the masking ratio ($m_{r}$) exceeds 0.5. The performance of the previously established TALL-Swin [53] also declines when the masking ratio surpasses 0.2. A similar trend is discernible in Fig. 3(c), which evaluates the DFDC testing set. The performance of contemporary methods diminishes at higher masking ratios, whereas the proposed GRACE method maintains relatively high performance even at a masking ratio of 0.8. We also draw the AUC comparison between the proposed GRACE and other peer methods in Fig. 3. We show that the proposed GRACE significantly outperforms other state-of-the-art DeepFake detectors, especially under noisy face sequences.
@@
3.3 Hyperparameters Selection
TABLE IV: Performance evaluation of the proposed GRACE with different hyperparameter settings using FF++ [1]. $g_{\text{dim}}$ and $g_{n}$ are the embedding dimension and number of layers of GCN, respectively; $N$ is the frames extracted from the video; $n_{\text{out}}$ is the number of neurons of FC; $\alpha$ stands for weights of sparsity.
$m_{r}$ $N$ ACC F1 AUC $g_{n}$ ACC F1 AUC
0.8 12 0.922 0.875 0.971 12 0.856 0.711 0.950
0.8 20 0.948 0.918 0.978 4 0.948 0.918 0.982
0.8 16 0.944 0.916 0.983 8 0.944 0.916 0.983
0.7 12 0.958 0.936 0.983 12 0.896 0.816 0.964
0.7 20 0.954 0.930 0.986 4 0.952 0.926 0.983
0.7 16 0.960 0.938 0.985 8 0.960 0.938 0.985
- $\alpha$ ACC F1 AUC $g_{\text{dim}}$ ACC F1 AUC
0.8 $1e^{-7}$ 0.928 0.887 0.978 600 0.896 0.805 0.966
0.8 $1e^{-6}$ 0.942 0.910 0.974 200 0.934 0.886 0.981
0.8 $1e^{-5}$ 0.944 0.916 0.983 400 0.944 0.916 0.983
0.7 $1e^{-7}$ 0.954 0.930 0.981 600 0.938 0.896 0.984
0.7 $1e^{-6}$ 0.944 0.914 0.979 200 0.942 0.905 0.987
0.7 $1e^{-5}$ 0.960 0.938 0.985 400 0.960 0.938 0.985
- $n_{\text{out}}$ ACC F1 AUC
0.8 1024 0.926 0.871 0.969
0.8 3072 0.924 0.876 0.939
0.8 2048 0.944 0.916 0.983
0.7 1024 0.944 0.908 0.984
0.7 3072 0.926 0.880 0.968
0.7 2048 0.960 0.938 0.985
To achieve optimal performance and robustness, we conducted a comprehensive ablation study to investigate the impact of various hyperparameters on the proposed GRACE method. This analysis provides valuable insights into the design choices and trade-offs involved in developing an effective DeepFake video detection system for real-world scenarios with noisy face sequences. Table IV presents the performance comparison of GRACE under different hyperparameter settings, evaluated on the challenging FF++ dataset [1].
3.3.1 Number of Extracted Frames ($N$)
The number of frames employed during the training and testing phases is a crucial aspect of GRACE. While using a larger number of frames might intuitively improve performance, it also significantly increases the computational complexity. To strike an optimal balance, we investigated the impact of varying the number of extracted frames. As shown in Table IV, using $N=8$ frames results in the lowest computational complexity but slightly compromises performance in terms of Accuracy, Macro F1-Score, and AUC. Conversely, increasing the number of frames to $N=20$ achieves state-of-the-art performance for most masking ratios during testing. Considering the trade-off between effectiveness and efficiency, we recommend using $N=16$ frames as the optimal setting for GRACE.
3.3.2 Number of GCN Layers ($g_{n}$)
The depth of the Graph Convolutional Network (GCN) plays a vital role in learning robust feature representations. However, stacking too many layers with the Graph Laplacian smooth prior may lead to over-smoothing of nodes and reduce the discriminative power. We explored the impact of varying the number of GCN layers ($g_{n}$) in GRACE. As presented in Table IV, setting $g_{n}=12$ results in suboptimal performance compared to $g_{n}=8$ and $g_{n}=4$, likely due to convergence difficulties within the given 200 epochs. While $g_{n}=4$ achieves outstanding performance overall, it slightly underperforms in highly noisy conditions (i.e., $m_{r}=0.8$) compared to $g_{n}=8$. Therefore, we suggest using $g_{n}=8$ as a balanced choice for stable and robust performance across various noise levels.
3.3.3 Sparsity Penalty Term ($\alpha$)
The sparsity penalty term $\alpha$ in the proposed loss function controls the balance between the sparsity constraint and the classification objective. A higher value of $\alpha$ encourages GRACE to learn a sparser feature representation, which is particularly beneficial for DeepFake video detection in the presence of invalid facial images. We investigated the impact of $\alpha$ by varying its value from $1e^{-7}$ to $1e^{-5}$. As shown in Table IV, a higher sparsity penalty enhances the network’s ability to learn essential and discriminative features, thereby reducing the influence of invalid faces and improving overall performance. However, setting $\alpha$ higher than $1e^{-5}$ leads to convergence difficulties. Based on our analysis, we recommend using $\alpha=1e^{-5}$ to achieve a balanced trade-off between sparsity and convergence stability.
3.3.4 GCN Embedding Dimension ($g_{\text{dim}}$)
The embedding dimension of the GCN ($g_{\text{dim}}$) determines the richness of the learned feature representations for DeepFake video detection. We investigated the impact of $g_{\text{dim}}$ by comparing the performance of GRACE with $g_{\text{dim}}\in{200,400,600}$, as shown in Table IV. Since the dimension of the graph representation ${\bm{A}}$ is $400\times 400$, intuitively, the best performance is achieved when $g_{\text{dim}}=400$. Reducing $g_{\text{dim}}$ below this value limits the expressive power of the GCN, while increasing it beyond introduces redundancy and harms performance. Therefore, we suggest setting $g_{\text{dim}}=400$ for optimal results.
3.3.5 Number of Fully Connected Layer Neurons ($n_{\text{out}}$)
To aggregate the output of the GCN and feed it into the softmax classifier, a simple fully connected (FC) layer is employed, projecting the graph representation to an $n_{\text{out}}$-dimensional feature vector. We investigated the impact of $n_{\text{out}}$ by comparing the performance of GRACE with $n_{\text{out}}\in{1024,2048,3072}$, as shown in Table IV. While $n_{\text{out}}=2048$ achieves excellent performance under highly noisy face sequences, the performance gap between $n_{\text{out}}=2048$ and $n_{\text{out}}=1024$ is insignificant, suggesting that the choice of $n_{\text{out}}$ is not highly sensitive. Based on our analysis, we recommend setting $n_{\text{out}}=2048$ for a good balance between performance and computational complexity.
@@ 3.4 Ablation Study
Table V presents an ablation study for the proposed modules in our GRACE, i.e., GCN, Graph Laplacian smooth prior, and Sparsity regularizer, where the performance is evaluated in noisy face sequences (say, $m_{r}=0.8$ and $m_{r}=0.7$). Note that when none of the proposed modules is adopted, we adopt the Transformer [46] as the classification head with four-head multi-head self-attention (MHSA) with the embedding size of 512 to meet a similar number of parameters with that of our GRACE, which could be treated as a variant of Convolutional Transformer. When we enable the GCN for the proposed feature entanglement and its affinity matrix, the performance of the DeepFake video detection under noisy face sequence, implying that the feature entanglement and its graph representation judiciously embeds the different spatiotemporal features into every node, thereby reducing the impact of invalid faces under noisy face sequences. Furthermore, the Graph Laplacian smooth prior could improve the robustness since it could filter noisy nodes that might contain many invalid faces without significantly increasing computational complexity. As shown in Fig.4, the convergence of the proposed GRACE with Graph Laplacian remains stable and shows outstanding performance on the FF++ [1] validation set with $m_{r}=0$. Finally, the sparsity regularizer significantly benefits the robustness of the DeepFake detection since we could enforce our GRACE learning essential and sparse features for the valid faces.
TABLE V: Ablation study of the proposed GRACE using different classification heads and components. GCN, GL, and Spa. denote the Graph Convolutional Network, Graph Laplacian smooth prior, and Sparsity regularizer, respectively.
$m_{r}$ GCN GL Spa. ACC F1 AUC #Param / MACs
0.8 0.844 0.655 0.705 40.79M, 38.82
@@
Figure 4: The validation accuracy curve evaluated on FF++ [1] of the proposed GRACE with and without Graph Laplacian and Sparsity regularizer.
@@ 3.5 Adversarial Attack on Face Detector
As previously discussed, DeepFake videos can be intentionally perturbed to evade detection by face detectors, rendering DeepFake detection ineffective. To emulate this real-world challenge, we employ an open-source adversarial attack on the MTCNN face detector [54], leveraging a PGD-like algorithm [39] with a maximum perturbation value of $\epsilon=0.04$, step size $\alpha_{\text{adv}}=0.01$, and the number of iterations $s=10$ to perturb the test sets of FF++ [1]. The step size $\alpha_{\text{adv}}$ determines the magnitude of each perturbation step applied to the input image during the iterative adversarial attack process. Assuming that each frame must contain at least one face detectable by MTCNN to assess the rate of missed detections, a black image will replace the frame when no face is detected. The results reveal that the average number of missed detected faces is 3.58 for the FF++ [1] test set, closely mirroring an $m_{r}=0.2$ scenario. However, the perturbed face sequences not only make it more difficult to recognize whether the video is fake due to the adversarial noise affecting the image quality, as exemplified in Fig. 5, but also introduce another challenge: the adversarial attack could cause the face detector to extract non-facial regions (e.g., background). This implies that the actual $m_{r}$ in this case could be even higher than the estimated $0.2$, as some of the detected faces might not be genuine facial regions. Despite these challenging conditions, our GRACE method maintains strong performance compared to other state-of-the-art methods, as illustrated in Table VI. Remarkably, the performance of all peer methods dropped significantly, partly due to the missed detections and partly because the adversarial noise introduces spatial distortions in the facial images. This real-world simulation further substantiates GRACE as a generalized, robust, and effective DeepFake detection model capable of handling noisy face sequences.
TABLE VI: The performance comparison of the proposed GRACE and other methods trained on FF++ under simulated real-world scenarios (i.e., adversarial attack on face detector).

### v3

@@ 4.2 Quantitative Results
TABLE II: Quantitative comparison of the noisy face sequences under different masking rations $m_{r}$ between the proposed GRACE and other state-of-the-art methods. We highlight the best performance in red and the second-best performance in blue, considering the several benchmark dataset, such as FF++ [1], DFDC [2], and Celeb-DF [3] with different $m_{r}$.
Method Venue Type $m_{r}$ FF++ [1] Celeb-DF [3] DFDC [2]
Acc. F1 AUC Acc. F1 AUC Acc. F1 AUC
Xception [6] CVPR2017 Image-based 0.0 0.925 0.894 0.972 0.861 0.806 0.910 0.953 0.910 0.981
@@
CORE [37] CVPRW2022 Image-based 0.0 0.948 0.925 0.984 0.953 0.940 0.989 0.950 0.903 0.977
0.4 0.883 0.799 0.888 0.858 0.790 0.890 0.907 0.781 0.870
0.8 0.818 0.601 0.663 0.764 0.572 0.661 0.863 0.595 0.651
UCF [41] CVPR2023 Image-based 0.0 0.937 0.911 0.982 0.856 0.792 0.891 0.890 0.815 0.939
@@
DFIL [39] ACMMM2023 Image-based 0.0 0.954 0.939 0.987 0.957 0.954 0.964 0.940 0.881 0.955
0.4 0.876 0.808 0.893 0.695 0.684 0.825 0.886 0.720 0.813
0.8 0.759 0.603 0.665 0.518 0.350 0.644 0.855 0.565 0.621
TALL-Swin [40] ICCV2023 Video-based 0.0 0.913 0.868 0.881 0.913 0.933 0.924 0.911 0.812 0.984
@@
DFGaze [42] TIFS2024 Video-based 0.0 0.946 0.926 0.986 0.956 0.954 0.972 0.915 0.881 0.968
0.4 0.854 0.724 0.795 0.756 0.726 0.824 0.818 0.709 0.743
0.8 0.785 0.652 0.656 0.612 0.669 0.659 0.798 0.648 0.596
GRACE [Ours] - Video-based 0.0 0.962 0.942 0.989 0.989 0.968 0.998 0.969 0.942 0.988
@@
TABLE III: Complexity comparison of different methods in terms of FLOPs , MACs, and #Params.
Method FLOPs (T) MACs (T) #Params (M)
@@
DFGaze [42] 22.647 11.241 123.185
GRACE [Ours] 70.751 35.246 29.661
@@
The primary performance assessment comparing the handling of invalid facial images between our proposed model, GRACE, and various state-of-the-art schemes is provided in Table II. Under optimal conditions, where most facial images are valid, GRACE exhibits competitive results, holding its own against other cutting-edge DeepFake video detection methods such as Xception [6], MesoNet [5], $F^{3}$-Net [43], RECCE [38], CORE [37], DFIL [39], TALL-Swin [40], and DFGaze [42]. It is worth noting that TALL-Swin and DFGaze are a video-based approaches.
@@
Specifically, the F1-Score of GRACE slightly surpasses those of its contemporaries under clean cases (i.e., $m_{r}=0$). This outcome implies that the proposed GRACE with Graph Laplacian Smoothing Prior Regularization is effective and reliable for DeepFake video detection. However, in scenarios where partial face images are invalid due to purposeful attacks on face detectors, the performance of traditional image-based methods, including Xception [6], MesoNet [5], $F^{3}$-Net [43], RECCE [38], CORE [37], UCF [41], and DFIL [39], may substantially deteriorate since they fail to consider noisy face sequences in real-world scenarios.
@@
Similarly, the video-level DeepFake detection methods, TALL-Swin [40] and DFGaze [42], which heavily rely on temporal cues, may suffer further performance degradation when the masking ratio increases. Invalid faces can cause landmark detection failures and incorrect temporal trajectories. Consequently, the F1-Score of TALL-Swin and DFGaze under a masking ratio of 0.8 in the testing phase are lower than 0.7, implying that all predictions would be categorized as either entirely fake or real. Likewise, the performance of another state-of-the-art video-based DeepFake detection method, TALL-Swin [40] and DFGaze [42], are poor when $m_{r}$ is increased. In stark contrast, all quality indices of our proposed GRACE, evaluated on different datasets, display promising results, suggesting that GRACE is robust and reliable even under highly noisy face sequences (e.g., when $m_{r}=0.8$). Remarkably, since most DeepFake detection methods fail to discuss the impact of unreliable face sequences, the degraded performance is most likely predictable.
@@
To further demonstrate the efficiency and practicality of the proposed method, we conduct a comprehensive complexity analysis and compare it with other state-of-the-art DeepFake detection methods. Table III presents the comparison results in terms of floating-point operations (FLOPs), multiply-accumulate operations (MACs), and the number of model parameters for each method with $16\times 3\times 144\times 144$ tensor for the fair comparison. It is evident that GRACE achieves a remarkable balance between computational complexity and performance. With 70.751 trillion FLOPs, 35.246 trillion MACs, and 29.661 million parameters, GRACE exhibits a moderate computational overhead compared to other methods, such as TALL-Swin [20], DFGaze [42], UCF [41], and RECCE [38]. Notably, GRACE outperforms these methods in terms of FLOPs and MACs while maintaining a comparable number of parameters. Moreover, GRACE demonstrates superior performance in handling noisy face sequences, as shown in the experimental results, despite having a similar complexity to methods like CORE [37], Xception [6], and DFIL [39]. This highlights the effectiveness of the proposed GRACE in learning discriminative and robust representations for DeepFake detection. The complexity analysis further substantiates GRACE as a practical and efficient solution for real-world DeepFake detection challenges, offering a compelling trade-off between computational resources and detection accuracy.
@@
Figure 3: The performance comparison of the proposed GRACE and other state-of-the-art methods in terms of AUC under different masking ratios $m_{r}$ for (a) FF++[1], (b) DFDC [2], and (c) Celeb-DF [3].
The detailed quantitative results, evaluated on benchmark datasets, are illustrated in Figures 3(a), 3(b) and 3(c). In the clean case, i.e., when $m_{r}=0$, the performance of the proposed method is comparable to other state-of-the-art methods. It is observed that performance degradation becomes increasingly pronounced with a rise in the masking ratio during the testing phase, particularly when the masking ratio ($m_{r}$) exceeds 0.5. The performance of the previously established TALL-Swin [40] and DFGaze [42] also decline when the masking ratio surpasses 0.2. A similar trend is discernible in Figure 3(c), which evaluates the DFDC testing set. The performance of contemporary methods diminishes at higher masking ratios, whereas the proposed GRACE method maintains relatively high performance even at a masking ratio of 0.8. We also draw the AUC comparison between the proposed GRACE and other peer methods in Figure 3. We show that the proposed GRACE significantly outperforms other state-of-the-art DeepFake detectors, especially under noisy and unreliable face sequences.
@@ 4.3 Ablation Study
Table IV presents an ablation study for the proposed modules in our GRACE, i.e., GCN, GLSPR, and SC, where the performance is evaluated in noisy face sequences (say, $m_{r}=0.8$ and $m_{r}=0.7$). Note that when none of the proposed modules is adopted, we adopt the Transformer [49] as the classification head with four-head multi-head self-attention (MHSA) with the embedding size of 512 to meet a similar number of parameters with that of our GRACE, which could be treated as a variant of Convolutional Transformer. When we enable the GCN for the proposed feature entanglement and its affinity matrix, the performance of the DeepFake video detection under noisy face sequence, implying that the feature entanglement and its graph representation judiciously embeds the different spatiotemporal features into every node, thereby reducing the impact of invalid faces under noisy face sequences. Furthermore, the proposed GLSPR could improve the robustness since it could filter noisy nodes that might contain many invalid faces without significantly increasing computational complexity. As shown in Figure 4, the convergence of the proposed GRACE with GLSPR remains stable and shows outstanding performance on the FF++ [1] validation set with $m_{r}=0$, demonstrating that the proposed GRACE still remains training smoothness and increased stably. Moreover, the proposed GRACE shows the better performance on the noisy face sequences, making it more practical. Finally, the proposed SC significantly benefits the robustness of the DeepFake detection since we could enforce our work learning essential and sparse features for the valid faces.
TABLE IV: Ablation study of the proposed GRACE using different classification heads and components.
$m_{r}$ GCN GLSPR SC Acc. F1 AUC #Param / MACs
0.8 0.844 0.655 0.705 40.79M / 38.82
@@
Figure 4: The validation accuracy curve evaluated on FF++ [1] of the proposed GRACE with and without Graph Laplacian Smoothing Prior Regularization and Sparsity Constraint.
@@ 4.4 Adversarial Attack on Face Detector
As previously discussed, DeepFake videos can be intentionally perturbed to evade detection by face detectors, rendering DeepFake detection ineffective. To emulate this real-world challenge, we employ an open-source adversarial attack on the MTCNN face detector [56], leveraging a PGD-like algorithm [17] with a maximum perturbation value of $\epsilon=0.04$, step size $\alpha_{\text{adv}}=0.01$, and the number of iterations $s=10$ to perturb the test sets of FF++ [1]. The step size $\alpha_{\text{adv}}$ determines the magnitude of each perturbation step applied to the input image during the iterative adversarial attack process. Assuming that each frame must contain at least one face detectable by MTCNN to assess the rate of missed detections, a black image will replace the frame when no face is detected. The results reveal that the average number of missed detected faces is 3.58 for the FF++ [1] test set, closely mirroring an $m_{r}=0.2$ scenario. However, the perturbed face sequences not only make it more difficult to recognize whether the video is fake due to the adversarial noise affecting the image quality, as exemplified in Figure 5, but also introduce another challenge: the adversarial attack could cause the face detector to extract non-facial regions (e.g., background). This implies that the actual $m_{r}$ in this case could be even higher than the estimated $0.2$, as some of the detected faces might not be genuine facial regions. Despite these challenging conditions, our method maintains strong performance compared to other state-of-the-art methods, as illustrated in Table V. Remarkably, the performance of all peer methods dropped significantly, partly due to the missed detections and partly because the adversarial noise introduces spatial distortions in the facial images. This real-world simulation further substantiates GRACE as a generalized, robust, and effective DeepFake detection model capable of handling noisy face sequences.
TABLE V: The performance comparison of the proposed GRACE and other methods trained on FF++ under simulated real-world scenarios (i.e., adversarial attack on face detector).